Please help me solve: sin3A=1, o<=A<2pi.
Thank you, Ray

Answer by Alan3354(21605)

Please help me solve: sin3A=1, o<=A<2pi.
------------
sin(3A) = 1
sin(90) = 1 [and sin(450), sin(810), every 360 degrees]
3A = 90, 450, 810
A = 30, 150, 270 within the specified limits
A = PI/6, 5*PI/6, 9*PI/6 (=3*PI/2) in radians